OpenStudy (anonymous):

You have a bag of chocolate candy. Ten of them are red, 7 are brown, 12 are green, and 9 are blue. What is the probability that you pick a red candy, given that you have already picked a red candy and have not replaced it? A) 10/37 B) 14/19 C) 5/19 D) 9/37 21) Which expression is equivalent to m4/2r3/2 in simplified form? A) mr B) m2 r C) m4 r3 D) m2 r r 22) For the function f(x) = -2x2, if the domain is {-3, 0, 3}, find the range. A) {0, -36} B) {0, -18} C) {-18, 0, 18) D) {-36, 0, 36} @Lily2913 @misty1212

Nnesha (nnesha):

can i get one chocolate ??

OpenStudy (anonymous):

no

Nnesha (nnesha):

(╯︵╰,)

Nnesha (nnesha):

what's the total number of chocolates ?

OpenStudy (anonymous):

lily

Nnesha (nnesha):

\[\large\rm \frac{ number ~of~red~chocolates }{total~number~of~chocolates }\] remember there are total 10 are red and we have already picked one red

OpenStudy (anonymous):

the total number of chocolates in the bag before you take any out is 38. since you the problem says that you've removed 1 red you now have 9 red and a total of 38-1 in the bag. so 9/37

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

these and two more questions lily

OpenStudy (anonymous):

for #22 For the function f(x) = -2x2, if the domain is {-3, 0, 3}, find the range. since f(x) = y and domain is = x you can find the range by doing -2(-3)2, -2*0*2, -2*3*2

OpenStudy (anonymous):

sorry I just saw that you meant it as -2x^2 ... anyway just plug in the values you have for x into -2x^2 and you'll get your range

OpenStudy (anonymous):

(-18,0,-18) or (0,-18)

OpenStudy (anonymous):

for 21) is m4/2r3/2 the same as \[m\frac{ 4 }{ 2 }r \frac{ 3 }{ 2 }\] ? if not, then I am not sure how to read it.

Nnesha (nnesha):

no offense plz. it was so disrespectful how you provided the answer when i asked question. http://prntscr.com/8xbbc1 link http://openstudy.com/code-of-conduct

Nnesha (nnesha):

2ndly we are not here to do their work but to guide them you are really good at explaining no doubt. i think it would be great if you let them do some work like adding/plugging in numbers instead doing on your own.